# Pose Anything: A Graph-Based Approach for Category-Agnostic Pose Estimation

### v1

@@ 4 Experiments
Support GT POMNet [42] CapeFromer [32] Baseline-T Ours-T
@@
Figure 5: Qualitative Results. We visualize the keypoint predictions under a 1-shot setting. The left column denotes the support image with its corresponding skeleton. The second column is the ground-truth query keypoints. The following columns are results from POMNet, CapeFormer, the enhanced baseline, and our method.
We use the MP-100 dataset [42] as our training and evaluation dataset, comprising samples drawn from existing category-specific pose estimation datasets. The MP-100 dataset encompasses a collection of over 20,000 images, spanning 100 distinct categories, with varying keypoint numbers up to 68 across different categories. To facilitate model training and evaluation, the samples are divided into five distinct splits. Notably, each split ensures that training, validation, and test categories are mutually exclusive, meaning that the categories used for evaluation remain unseen during the training process.
The dataset includes partial skeleton annotations in various formats, including differences in the indexing of keypoints (some start from zero while others from one). To standardize and enhance the dataset, we adopted a unified format with comprehensive skeleton definitions for all categories. This process involved cross-referencing the original datasets and, where necessary, conducting manual annotations to ensure completeness and consistency.
To quantify our model’s performance, we employ the Probability of Correct Keypoint (PCK) metric [47], with a PCK threshold set at 0.2, following the conventions established by POMNet [42] and CapeFormer [32].
@@
We show a qualitative comparison between our method and previous CAPE methods CapeFormer [32] and POMNet [^1] We used POMNet official code to train and visualize results, as they don’t provide pre-trained models. Quantitative results are taken from the published paper. [42] in Figure 5. The last column denotes our full method, which is based on our enhanced baseline, replacing the previous decoder with our novel GTD module. As can be seen, the enhanced baseline achieves better keypoint localization than the previous state-of-the-art. Furthermore, the structure information incorporated in our method serves as a strong prior for localizing keypoints, helping in breaking symmetry and creating structure consistency between the keypoints. The first row of the figure shows an example of symmetry that confuses prior methods but not ours, while the third row shows our method’s superiority in preserving structure consistency across keypoints. More examples are in the supplementary.
4.1 Implementation Details
For a fair comparison, network parameters, training parameters, and data augmentations and pre-processing are kept the same as CapeFormer [32]: Both the encoder and decoder have three layers and $\lambda_{heatmap}$ is set to $2.0$. The model is built upon MMPose framework [6], trained using Adam optimizer for 200 epochs with a batch size of 16, the learning rate is $10^{-5}$, and decays by 10× on the 160th and 180th epoch. More design choices and evaluations are in the supplementary.
Enhanced Backbone.
We use a stronger, transformer-based, SwinV2-T [22] backbone, instead of the original ResNet-50 [14] backbone, which can also provide a multi-scale feature map. After experimenting with various configurations, including multi-scale and single-scale feature maps, we found that applying a straightforward bilinear upsampling on the final feature layer yielded similar results while preserving simplicity. Additionally, we optimized the extraction of support keypoint features by employing a Gaussian kernel mask with lower variance, taking advantage of the enhanced feature quality of the new backbone. These simple adjustments led to an improvement of 3.2%.
Disable Support Keypoint Identifier.
CapeFormer introduced a keypoint positional encoding, termed the ”Support Keypoint Identifier”. This encoding is generated via sinusoidal encoding of the keypoints’ order. This encoding improves performance by a significant 3% PCK, at the cost of introducing a dependency on the order of keypoints. Although the dataset contains diverse object categories, some of them share a common keypoint ordering.
This became evident when we evaluated a model trained with this positional encoding on data of keypoints in reverse order. The result was a notable decrease of approximately 30% in PCK, indicating that the model had become highly specialized for a specific keypoint format. This sharp decrease is not present when the positional encoding is not used. We believe that category-agnostic pose estimation (CAPE) should not rely on such assumptions and should accommodate support keypoints without imposing a specific order. Therefore, we have chosen to remove this encoding from our baseline.
4.2 Benchmark Results
Model 1-Shot 5-Shot
Split 1 Split 2 Split 3 Split 4 Split 5 Avg Split 1 Split 2 Split 3 Split 4 Split 5 Avg
ProtoNet [33] 46.05 40.84 49.13 43.34 44.54 44.78 60.31 53.51 61.92 58.44 58.61 58.56
MAML [10] 68.14 54.72 64.19 63.24 57.20 61.50 70.03 55.98 63.21 64.79 58.47 62.50
Fine-tuned [26] 70.60 57.04 66.06 65.00 59.20 63.58 71.67 57.84 66.76 66.53 60.24 64.61
POMNet [42] 84.23 78.25 78.17 78.68 79.17 79.70 84.72 79.61 78.00 80.38 80.85 80.71
CapeFormer [32] 89.45 84.88 83.59 83.53 85.09 85.31 91.94 88.92 89.40 88.01 88.25 89.30
Baseline-T 89.48 86.69 85.31 84.79 84.97 86.25 94.04 91.20 89.19 90.63 89.45 90.90
Ours-T 91.06 88.24 86.09 86.17 85.78 87.47 94.18 91.46 90.50 90.18 89.47 91.12
Baseline-S 92.88 89.11 89.16 87.19 88.73 89.41 96.50 92.62 91.68 91.23 91.28 92.66
Ours-S 93.66 90.42 89.79 88.68 89.61 90.43 96.51 92.15 91.99 92.01 92.36 93.00
@@ 4.2 Benchmark Results
Table 1: MP-100 Results. PCK performance under 1-shot and 5-shot settings. In both settings, our approach consistently outperforms other methods, achieving the scores across all splits. *-T denotes using the Tiny version of the backbone and *-S using the Small version.
@@
We compare our method with the previous CAPE methods CapeFormer [32] and POMNet [42] and three baselines: ProtoNet [33], MAML [10], and Fine-tuned [26]. Further details on these model’s evaluation can be found in [42].
We report results on the MP-100 dataset under 1-shot and 5-shot settings in Table 1. As can be seen, the enhanced baseline models, which are agnostic to the keypoints order as opposed to CapeFormer, outperform previous methods and improve the average PCK by 0.94% under the 1-shot setting and 1.60% under the 5-shot setting. Our graph-based method further improves performance, improving the enhanced baseline by 1.22% under the 1-shot setting and 0.22% under the 5-shot setting, achieving new state-of-the-art results for both settings.
@@
We also show the scalability of our design. Similar to DETR-based models, employing a larger backbone improves performance. We show that our graph decoder design also enhances the performance of the larger enhanced baseline, improving results by 1.02% and 0.34% under 1-shot and 5-shot settings respectively.
Out-of-Distribution Performance.
To assess the robustness of our model, we evaluate the small version of our network using images from different domains. Results are shown in Figure 6. Our model, which was trained on real images only, demonstrates its adaptability and effectiveness across varying data sources such as cartoons and imaginary animals, created using a diffusion model. Furthermore, our model demonstrates satisfactory performance even when the support and query images are from different domains.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@ 4.3 Ablation Study
We conducted a number of ablation studies on the MP-100 dataset. We first evaluate our method using different backbones, showing the advantages of Swin Transformer architecture for localization tasks. We then show the contribution of the geometrical structure prior by evaluating our model using wrong skeleton relations. Lastly, we further show the power of graph structure by evaluating the performance using masked inputs. We perform all ablation experiments on the test set of MP-100 split1 under the 1-shot setting following [32, 42].
[…]
ResNet-50 [14] 25.5 4.1 87.71
DinoV1-S [4] 22.1 46.7 89.17
DinoV2-S [28] 22.1 46.7 90.34
SwinV2-T [22] 28.35 4.4 91.06
SwinV2-S [22] 49.73 8.57 93.66
Table 2: Pre-Trained Backbone Ablation. PCK scores using different pre-trained backbones. Swin transformer combines superior results with efficient processing.
Different Backbones.
We assess our model’s performance with various pre-trained backbones, including a CNN-based backbone (ResNet-50) and two distinct pre-trained transformer backbones, namely Dino [4, 28] and Swin V2 [22]. Dino and DinoV2 are self-supervised Vision Transformers trained via self-distillation, resulting in robust, finely detailed semantic features. Moreover, these encoded semantic representations are shared among related but distinct object categories. Swin Transformer restructures self-attention with a window mechanism, balancing efficiency and performance for vision tasks. This hierarchical architecture has the flexibility to model at different scales and has linear computational complexity with respect to image size. As shown in Table 2, SwinV2 outperforms other options, delivering superior results while maintaining efficiency comparable to the CNN-based backbone. In addition, using a larger backbone boosts performance, but at the price of efficiency and size.
The Contribution of Graph Structure.
To assess the contribution of the graph structure prior, we evaluate our method using random graph inputs by choosing edge connections randomly for each instance. This results in a performance decrease of 9.57%, validating the contribution of the graph decoder and the use of structural knowledge.
To highlight the power of graph information, we applied partial masking to the support/query image before executing our algorithm. It’s important to note that the difference between the baseline model and ours lies in the novel graph-based feed-forward network. Quantitative analysis in Figure 6(a) shows our method consistently outperforming the baseline. The dashed horizontal line represents the identity operation which outputs the support keypoint’s locations, preserving structure. Notably, our model can predict keypoints even when a significant portion of the support image is masked. This suggests that the model has learned which keypoints are relevant for each category and matches them to support features based on structure. For instance, in Figure 6(b) (top), the model accurately predicts the sofa’s structure and most keypoints. However, when large portions of the query image are masked, the model’s performance rapidly declines, although it retains structure, as evident in Figure 6(b) (bottom). More examples, and additional ablation studies, can be found in the supplementary.
@@
Support GT Baseline Ours
Figure 7: Masking. (a) Quantitative results when masking the support/query images. Our method consistently surpasses the baseline, leveraging cues provided by the graph structure to overcome information gaps in the support/query images. (b) Qualitative results when masking the support (top) and query (bottom) image using the enhanced baseline and our graph-based model.
@@ 5 Conclusion
We present a novel approach for category-agnostic pose estimation (CAPE) by recognizing the importance of underlying geometrical structures within objects. We introduce a robust Graph Transformer Decoder that significantly enhances the accuracy of keypoint localization by capturing and incorporating structural information, thus exploiting the relationships and dependencies between keypoints. In addition, we provide an updated version of the MP-100 dataset, which now includes skeleton annotations for all categories, further promoting research in CAPE.
Our experimental results demonstrate the superiority of our method over the previous state-of-the-art approach, CapeFormer, by a significant margin. With improvements under both 1-shot and 5-shot settings, our method showcases its scalability and efficiency opening the door to more versatile and adaptable applications in computer vision.
6 Acknowledgments
This work was partly funded by the Weinstein Institute.

### v2

@@ 4 Experiments
Following previous CAPE works, we use the MP-100 dataset [37] for training and evaluation, comprising samples drawn from existing category-specific pose estimation datasets. The MP-100 dataset encompasses a collection of over 20,000 images, spanning 100 distinct categories, with varying keypoint numbers up to 68 across different categories. To ensure categories used for evaluation remain unseen during the training process, the samples are divided into five different mutually exclusive splits. The dataset includes partial skeleton annotations in various formats, including differences in the indexing of keypoints (some start from zero while others from one). We adopted a unified format with comprehensive skeleton definitions for all categories (skeleton annotations are shared among categories). This process involved cross-referencing the original datasets and, where necessary, conducting manual annotations.
To quantify our model’s performance, we employ the Probability of Correct Keypoint (PCK) metric [42], with a PCK threshold set at 0.2, following previous conventions [37, 29].
4.0.1 CapeFormer-T.
In the following sections, we compare our model to previous works along with an updated version of the previous state-of-the-art method which we call CapeFormer-T. For this baseline, we replaced the original ResNet-50 [13] backbone with a stronger transformer-based SwinV2-T [21]. After testing various configurations, including multi-scale and single-scale feature maps, we found that applying bilinear upsampling on the final feature layer achieved comparable results while maintaining simplicity. To capitalize on the improved feature quality we extract keypoint features using a Gaussian kernel mask with lower variance. We believe that CAPE should not enforce a specific keypoint order. Therefore, following the discussion in the previous section, we removed keypoint positional encoding. These simple adjustments led to an improvement of 0.94% over the original CapeFormer. We provide a short description of the different components of CapeFormer in the supplementary.
It is important to note that the difference between CapeFormer-T and our method GraphCape, is the graph-based decoder design, which utilizes the graph structure during training and inference. Thus any performance difference between these models is a result of using our graph-oriented architecture.
4.1 Implementation Details
For a fair comparison, network parameters, training parameters, data augmentations and data pre-processing are kept the same as CapeFormer [29]: Both the encoder and decoder have three layers and $\lambda_{heatmap}$ is set to $2.0$. The model is built upon MMPose framework [5], trained using Adam optimizer for 200 epochs with a batch size of 16, the learning rate is $10^{-5}$, and decays by 10× on the 160th and 180th epoch. More design choices and evaluations are in the supplementary.
4.2 Qualitative Results
Support GT POMNet CapeFromer CapeFormer-T GraphCape
@@ 4.2 Qualitative Results
Figure 5: Qualitative Results. We visualize the keypoint predictions under a 1-shot setting. The left column denotes the support image with its corresponding skeleton. The second column is the ground-truth query keypoints. The following columns are results from POMNet, CapeFormer, CapeFormer-T, and our method.
@@
We show a qualitative comparison between our method, CapeFormer-T, and previous CAPE methods: CapeFormer [29] and POMNet [^1] We used POMNet official code to train and visualize results, as they don’t provide pre-trained models. Quantitative results are taken from the published paper. [37] in Figure 5. As can be seen, the structure information incorporated in our method serves as a strong prior for localizing keypoints, helping in breaking symmetry and creating structure consistency between the keypoints. The first row of the figure shows an example of symmetry that confuses prior methods but not ours, while the third row shows our method’s superiority in preserving structure consistency across keypoints. More examples are in the supplementary.
4.3 Quantitative Results
@@
Table 2: MP-100 Results. PCK performance under 1-shot and 5-shot settings. In both settings, our approach consistently outperforms other methods, across all splits.
Model Split 1 Split 2 Split 3 Split 4 Split 5 Avg
1-Shot ProtoNet [30] 46.05 40.84 49.13 43.34 44.54 44.78
MAML [9] 68.14 54.72 64.19 63.24 57.20 61.50
Fine-tuned [25] 70.60 57.04 66.06 65.00 59.20 63.58
POMNet [37] 84.23 78.25 78.17 78.68 79.17 79.70
CapeFormer [29] 89.45 84.88 83.59 83.53 85.09 85.31
CapeFormer-T 89.48 86.69 85.31 84.79 84.97 86.25
GraphCape 91.19 87.81 85.68 85.87 85.61 87.23
5-Shot ProtoNet [30] 60.31 53.51 61.92 58.44 58.61 58.56
MAML [9] 70.03 55.98 63.21 64.79 58.47 62.50
Fine-tuned [25] 71.67 57.84 66.76 66.53 60.24 64.61
POMNet [37] 84.72 79.61 78.00 80.38 80.85 80.71
CapeFormer [29] 91.94 88.92 89.40 88.01 88.25 89.30
CapeFormer-T 94.04 91.20 89.19 90.63 89.45 90.90
GraphCape 94.24 91.32 90.15 90.37 89.73 91.16
@@
We compare our method with CapeFormer-T, previous CAPE methods CapeFormer [29] and POMNet [37], and three baselines: ProtoNet [30], MAML [9], and Fine-tuned [25]. Further details on these model’s evaluation can be found in [37]. We report results on the MP-100 dataset under 1-shot and 5-shot settings in Table 2. As can be seen, our graph-based method consistently improves performance over CapeFormer-T on the different dataset splits, on average by 0.98% under the 1-shot setting and 0.26% under the 5-shot setting, achieving new state-of-the-art results for both settings.
@@
In the supplementary, we also show better scalability of our design. Similar to DETR-based models, employing a larger backbone improves performance and our graph-based design also enhances performance using larger backbones. Furthermore, we show that our graph model scales better with the number of decoder layers compared to CapeFormer-T.
@@ 4.4 Ablation Study
We conducted several ablation studies on the MP-100 dataset. We first show quantitative examples using out-of-distribution images. Then, we display the contribution of the geometrical structure prior by evaluating our model using different skeleton relations. Then, we demonstrate the advantages of using graph structure in handling occlusions by evaluating the performance using masked inputs. Lastly, we test our model’s performance under a cross-category setting. We perform all ablation experiments on the test set of MP-100 split1 under the 1-shot setting following [29, 37]. In the supplementary, we also show our model’s performance with various pre-trained backbones, aligning with DETER-based results [19] regarding the preferred backbone.
4.4.1 Out-of-Distribution Performance.
To assess the robustness of our model, we evaluate our network using images from different domains. Results are shown in Figure 6. Our model, which was trained on real images only, demonstrates its adaptability and effectiveness across varying data sources such as cartoons and imaginary animals, created using a text-to-image diffusion model. Furthermore, our model performs satisfactorily even when the support and query images are from different domains. However, when the images are from vastly different categories from the ones seen during training, the performance degrades. Yet, this is common to all CAPE methods and doesn’t relate to using graphs.
@@
(a) (b)
4.4.2 The Contribution of Graph Structure.
To assess the contribution of the graph structure prior, we evaluate our method using different graph inputs. Results are shown in Table 3. Training and evaluating the graph model without node’s connectivity, the model is equivalent to CapeFormer-T, thus performing the same. Moreover, training the model on fully-connected graphs results in a similar performance to CapeFormer-T as mathematically it is equivalent to adding a bias vector to all nodes, which is normalized and canceled. And again, performance is similar to CapeFormer-T (intuitively, as all graph inputs are the same, no new information is provided). When providing random graphs the model’s performance degrades as wrong connectivity involves inferior semantic features for localization. This shows a limitation of using graphs, as the wrong skeleton data may impair performance. Lastly, we show in the supplementary that different graph definitions may result in different performances. However, we noticed empirically that most straightforward skeletal definitions produce similar results. Furthermore, as we didn’t optimize the dataset skeletons, we believe that further quantitative improvements can be made using optimal skeletal relations. Thus, we defer the exploration of the optimal skeletons to future research.
Table 3: Different Graph Connectivity: We evaluate our graph-oriented network, training and evaluating with different types of graph definitions.
Graph Structure PCK0.2
No Connectivity 89.48
Fully Connected 89.35
Random Graphs (each instance) 81.49
Random Graphs (each category) 88.10
Pose Graph 91.19
4.4.3 Masking Support/Query Images.
To highlight the contribution of graph information in handling occlusions, we applied partial masking to the support/query image before executing our algorithm. It is important to note that the difference between the two compared models is our suggested graph-based decoder. Quantitative analysis in Figure 7(a) shows our method consistently outperforming CapeFormer-T. The dashed horizontal line represents the identity operation which outputs the support keypoint’s locations, preserving structure. Notably, our model can predict keypoints even when a significant portion of the support image is masked. This suggests that the model has learned which keypoints are relevant and matches them based on structure. For instance, in Figure 7(b) (top), the model accurately predicts most keypoints. In comparison, CapeFormer-T is confused by symmetry, breaking the structure of the sofa. However, when large portions of the query image are masked, the model’s performance rapidly declines, although it retains structure, as evident in Figure 7(b) (bottom). This experiment suggests that our model better handles occlusions, which is crucial for real-world applications. More examples are in the supplementary.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Support GT CapeFormer-T GraphCape
Figure 7: Masking. (a) Quantitative results when masking the support/query images. Our method consistently surpasses CapeFormer-T, leveraging cues provided by the graph structure to overcome information gaps in the support/query images. (b) Qualitative results when masking the support (top) and query (bottom) image using CapeFormer-T and our graph-based model. Our model better handles occlusions by preserving structure and breaking symmetry.
4.4.4 Cross-Category Correspondence.
We test our model’s performance when the support and query images are from vastly different categories. Results can be seen in Figure 8. CapeFormer-T correctly finds correspondence between the legs of the chair and of the horse. However, it mistakenly matches keypoints according to their location in the image and not according to the orientation of the objects. Our model breaks the symmetry between the leg keypoints and correctly matches keypoints with the correct 3D orientation.
Refer to caption Refer to caption Refer to caption
Support Capefromer-T GraphCape
Figure 8: Cross-Category Correspondence. We test our model’s performance when the support and query images are from vastly different categories. The difference lies in matching keypoints. CapeFormer-T matches keypoints naively, based on their location in the image. In contrast, our model breaks the symmetry between the legs and correctly matches keypoints with the correct object 3D orientation.
@@ 5 Conclusion
We present a novel approach for category-agnostic pose estimation (CAPE) by treating input keypoints as graphs, recognizing the importance of underlying geometrical structures within objects. We implement this prior using a Graph-FFN, which captures and incorporates structural information by exploiting the relationships and dependencies between keypoints. This change significantly enhances the accuracy of keypoint localization. In addition, we provide an updated version of the MP-100 dataset, which now includes skeleton annotations for all categories, further promoting research in CAPE.
Our experimental results demonstrate the superiority of our method over our updated version of the previous state-of-the-art approach, CapeFormer-T. With improvements under both 1-shot and 5-shot settings, our method opens the door to more versatile and adaptable applications in computer vision.
5.0.1 Acknowledgement.
Part of this research was supported by the Weinstein Institute and ISF grant 2132/23.